### v1

# Learning Effective NeRFs and SDFs Representations with 3D Generative Adversarial Networks for 3D Object Generation: Technical Report for ICCV 2023 OmniObject3D Challenge

Abstract
In this technical report, we present a solution for 3D object generation of ICCV 2023 OmniObject3D Challenge. In recent years, 3D object generation has made great process and achieved promising results, but it remains a challenging task due to the difficulty of generating complex, textured and high-fidelity results. To resolve this problem, we study learning effective NeRFs and SDFs representations with 3D Generative Adversarial Networks (GANs) for 3D object generation. Specifically, inspired by recent works, we use the efficient geometry-aware 3D GANs as the backbone incorporating with label embedding and color mapping, which enables to train the model on different taxonomies simultaneously. Then, through a decoder, we aggregate the resulting features to generate Neural Radiance Fields (NeRFs) based representations for rendering high-fidelity synthetic images. Meanwhile, we optimize Signed Distance Functions (SDFs) to effectively represent objects with 3D meshes. Besides, we observe that this model can be effectively trained with only a few images of each object from a variety of classes, instead of using a great number of images per object or training one model per class. With this pipeline, we can optimize an effective model for 3D object generation. This solution is one of the final top-3-place solutions in the ICCV 2023 OmniObject3D Challenge.
Figure 1: An overview of the framework of our solution for 3D object generation.
1 Introduction
3D object generation aims at generating meaningful 3D surfaces and synthesis images of 3D objects given random inputs [23, 7]. Inspired by the great success of 2D object generation [9, 4], there have been many studies [12, 5] extending 2D methods to 3D generation. Recent 3D generation approaches focus more on effective 3D representation learning [18, 15] and advanced generation strategy [10, 2, 21]. Despite the great progress and promising results in recent years, 3D object generation is still a challenging task due to the difficulty of generating complex, textured and high-fidelity results.
To devise a solution for 3D object generation in the ICCV 2023 OmniObject3D Challenge, we simultaneously consider effective 3D representation learning and advanced generation strategy. In terms of 3D representation learning, textured mesh [3] and Neural Radiance Fields (NeRFs) [14] embrace the great potential to effectively encode 3D object features. In particular, textured mesh representations can support individual texture map generation, which allows us to easily replace surface textures, while NeRFs-based representations excel at rendering high-fidelity images. Furthermore, when considering effective shape representation, there are a variety of options, ranging from explicit representations like point-based [20], mesh-based [1] and voxel-based [22] approaches to implicit ones such as the Signed Distance Functions (SDFs) [17, 11] and occupancy function [13]. In the light of these cutting-edge techniques, we mainly focus on exploring effective NeRFs representations for 3D object generation while still optimizing SDFs representations to effectively represent objects with 3D meshes.
Moving on to advanced generation strategy, there are usually different options, such as latent generation [21] and direct generation [16]. In particular, latent generation methods aim to generate the representation of the desired output in the latent space and decode the latent code back to the 3D space, while direct generation focus on directly generating the desired 3D representation without utilizing the latent space. Besides, from the perspective of generation process, generation strategies can be categorized into Generative Adversarial Networks (GANs) [10, 2], Variational AutoEncoder (VAE) [19], flow-based [6], Diffusion Models (DMs) [21], etc. In our solution, we employ the efficient geometry-aware 3D GANs [2] as the backbone for 3D object generation.
The overall framework of our solution is depicted in Fig. 1. Specifically, we aim to learn effective NeRFs and SDFs representations with 3D Generative Adversarial Networks (GANs) for 3D object generation. Inspired by the recent success of efficient geometry-aware 3D GANs, i.e., EG3D [2, 18], we employ it as the backbone and incorporate label embedding and color mapping [18] to enable model training on different taxonomies simultaneously. Next, we generate NeRFs-based representations for rendering high-fidelity synthetic images and SDFs to effectively represent objects with 3D meshes through a decoder. In addition, we empirically find that we can effectively train this model with only a few images per object from various classes, rather than training with a great number of images per object or optimizing one model per class. With this pipeline, we can optimize an effective model for 3D object generation. This solution is one of the final top-3-place solutions in the ICCV 2023 OmniObject3D Challenge.
2 Technical Solution
As aforementioned, to realize 3D object generation for ICCV 2023 OmniObject3D Challenge, we explore 3D GANs to learn effective NeRFs representations for synthetic images and SDFs representations for 3D shapes.
Motivation of Using GANs instead of DMs.
Although diffusion models have recently gained an increasing popularity in 3D object generation, most existing works [16, 21] have only been verified to work under specific category conditions (e.g., chairs [16] or avatars [21]). Besides, since the OmniObject3D dataset [23] is a collection of objects from hundreds of taxonomies, following the pipeline of existing DMs would necessitate the use of hundreds of pre-trained models, which is impractical. Hence, we employ a method that enables simultaneous training on all the taxonomies in the OmniObject3D dataset. Moreover, given the success of StyleGAN2 [10] in representing NeRFs in latent space, we have chosen latent generation for our pipeline.
Solution Pipeline.
The framework of our solution is depicted in Fig. 1. Overall, our solution adopts a GAN-based generation process. Specifically, with EG3D [2] as the backbone, firstly, we generate a one-dimensional latent code, $\mathbf{z}$ ($1\times 512$), which is obtained by randomly sampling 512 values from a standard normal distribution. Meanwhile, we randomly generate an integer, indicating the object label, within the range of the total category number of OmniObject3D. The integer is then transformed into a label embedding by an embedding layer. This enables the model to be trained on different taxonomies simultaneously. Next, we forward the label embedding and latent code $\mathbf{z}$ into a mapping network and generate an intermediate latent code $\mathbf{w}$ ($1\times 15\times 512$) with promoted dimensions. Here, the first 14 dimensions of $\mathbf{w}$ ($\mathbf{w}_{1}$ in the figure) are used as the input to the object generator ($\mathbf{G}_{1}$ in the figure) while the last dimension of $\mathbf{w}$ ($\mathbf{w}_{2}$ in the figure) are employed as the input to the color generator ($\mathbf{G}_{2}$ in the figure). $\mathbf{G}_{1}$ and $\mathbf{G}_{2}$ generate the tri-plane representation ($3\times 32\times 64\times 64$) and the color tensor ($3\times 10$), respectively. Following [18], we use the color generator which allows convenient texture conversion. After that, we determine queries, $\mathbf{Q}$, based on the randomly generated camera pose. The tri-plane representation is then taken as keys, $\mathbf{K}$, and transmitted into the tri-plane decoder together with $\mathbf{Q}$. Then, we use a decoder to generate the SDF of the object, which can be sampled to create a 3D mesh. This decoder also produces a feature as the input to a view direction mapper, from which the output is multiplied by the color tensor, $\mathbf{V}$ (the values), in the attention module. Finally, the attention module produces the 2D synthesis image and a discriminator [2] which only functions during training is used for adversarial training.
Model Training.
Since we are simultaneous optimizing NeRFs and SDFs representations, we ensure the generator can initially generate a unit sphere by initializing our model with SDF-pretraining as [18]. Then, the generator is optimized using the Adam optimizer with a learning rate of 0.0025, while the discriminator [2] uses a learning rate of 0.002. The batch size is set to 32 and the model is trained for 300,000 iterations. Additionally, we also use adaptive discriminator augmentation [8] to improve model generalization. Overall, the model training objective is formulated as:
$\alpha_{0}\mathcal{L}_{path}+\alpha_{1}\mathcal{L}_{e}+\alpha_{2}\mathcal{L}_{v}+\alpha_{3}\mathcal{L}_{sdf},$ (1)
where $\mathcal{L}_{path}$ denotes the Path length regulation loss [10], $\mathcal{L}_{e}$ and $\mathcal{L}_{v}$ denote the entropy loss and total variation loss [10], respectively, $\mathcal{L}_{sdf}$ represents the SDF eikonal loss [18], and $\alpha_{i}$ are weighting parameters. We set $\alpha_{0}{=}2$, $\alpha_{1}{=}0.05$, $\alpha_{2}{=}0.5$ and $\alpha_{3}{=}0.1$.
Figure 2: Visualization results of the synthesis images rendered from the NeRFs generated by our solution.
3 Experiments
Dataset.
We train our model from scratch for 3D object generation on the training dataset provided by the ICCV 2023 OmniObject3D Challenge. The complete OmniObject3D dataset [23] contains 216 classes from a wide range of daily categories and approximately 6k objects in total where each object has 100 images in different views.
Efficient Model Training.
Instead of training one model for each class, we train a single model with various types of objects. The model takes each image along with its camera pose, focal length and object type. Besides, we empirically find that we can effectively train this model with only a few images per object from various classes, so we select the first 8 images of each object for model training. This significantly improves the model training efficiency.
Figure 3: Visualization results of the synthesis images rendered from the NeRFs generated by our solution without additional label embedding.
Method FID
$\Delta$= Ours w/ label embedding - -4.8
Ours w/o label embedding
Table 1: Improvement in terms of FID of our solution after using the additional label embedding.
Results.
Fig. 2 shows some visualization results of synthesis images generated by our solution on OmniObject3D. It can be seen that our approach is capable of generating high-fidelity images from different views for 3D object generation. Besides, Fig. 3 shows results of our solution without the additional label embedding. Comparing Fig. 2 with Fig. 3, we can see that without the additional label embedding, the quality of the synthesis images is significantly worse than the original solution. Furthermore, Table. 1 shows that using the additional label embedding can improve the solution by 4.86 in terms of FID. This verifies that the additional label embedding is beneficial for improving model generalization.
4 Conclusion
In this technical report, we present a solution for 3D object detection of ICCV 2023 OmniObject3D Challenge. The key insight of our solution is to learn effective NeRFs and SDFs Representations with 3D generative adversarial networks for 3D object generation. Overall, this solution is one of the final top-3-place solutions in the ICCV 2023 OmniObject3D Challenge.

### v2

@@
-Learning Effective NeRFs and SDFs Representations with 3D Generative Adversarial Networks for 3D Object Generation: Technical Report for ICCV 2023 OmniObject3D Challenge
+Learning Effective NeRFs and SDFs Representations with 3D GANs for Object Generation
 Abstract
-In this technical report, we present a solution for 3D object generation of ICCV 2023 OmniObject3D Challenge.
+We present a solution for 3D object generation of ICCV 2023 OmniObject3D Challenge.
 In recent years, 3D object generation has made great process and achieved promising results, but it remains a challenging task due to the difficulty of generating complex, textured and high-fidelity results.
@@
 With this pipeline, we can optimize an effective model for 3D object generation.
-This solution is one of the final top-3-place solutions in the ICCV 2023 OmniObject3D Challenge.
-Figure 1: An overview of the framework of our solution for 3D object generation.
-1 Introduction
+This solution is among the top 3 in the ICCV 2023 OmniObject3D Challenge.
+Index Terms:
+3D Object Generation, Generative Adversarial Networks, NeRFs, SDFs
+I Introduction
 3D object generation aims at generating meaningful 3D surfaces and synthesis images of 3D objects given random inputs [1, 2].
@@
 Furthermore, when considering effective shape representation, there are a variety of options, ranging from explicit representations like point-based [14], mesh-based [15] and voxel-based [16] approaches to implicit ones such as the Signed Distance Functions (SDFs) [17, 18] and occupancy function [19].
-In the light of these cutting-edge techniques, we mainly focus on exploring effective NeRFs representations for 3D object generation while still optimizing SDFs representations to effectively represent objects with 3D meshes.
+In light of these cutting-edge techniques, we mainly focus on exploring effective NeRFs representations for 3D object generation while still optimizing SDFs representations to effectively represent objects with 3D meshes.
+Figure 1: An overview of the framework of our solution for 3D object generation.
 Moving on to advanced generation strategy, there are usually different options, such as latent generation [11] and direct generation [20].
 In particular, latent generation methods aim to generate the representation of the desired output in the latent space and decode the latent code back to the 3D space, while direct generation focus on directly generating the desired 3D representation without utilizing the latent space.
-Besides, from the perspective of generation process, generation strategies can be categorized into Generative Adversarial Networks (GANs) [10, 2], Variational AutoEncoder (VAE) [19], flow-based [6], Diffusion Models (DMs) [21], etc.
+Besides, from the perspective of generation process, generation strategies can be categorized into Generative Adversarial Networks (GANs) [9, 10], Variational AutoEncoder (VAE) [21], flow-based [22], Diffusion Models (DMs) [11], etc..
 In our solution, we employ the efficient geometry-aware 3D GANs [10] as the backbone for 3D object generation.
@@
 This solution is one of the final top-3-place solutions in the ICCV 2023 OmniObject3D Challenge.
-2 Technical Solution
-As aforementioned, to realize 3D object generation for ICCV 2023 OmniObject3D Challenge, we explore 3D GANs to learn effective NeRFs representations for synthetic images and SDFs representations for 3D shapes.
-Motivation of Using GANs instead of DMs.
+Figure 2: Images rendered by our solution with label embedding.
+II Related Work
+3D assets [23, 1, 24] are crucial for robotics and computer vision [25] because they provide accurate spatial representations and realistic environments essential for training, simulation, and object recognition.
+Unfortunately, the collection of 3D assets is time-consuming and labor-intensive [26, 1].
+For example, when constructing the large-scale dataset, OmniObject3D [1], the objects are placed in a calibrated environment [1, 26] built using fiducial markers [27, 28, 29], and the 3D scanners sample the objects from different perspectives.
+These fiducial markers, though low-cost [30, 31, 32, 33, 34], require additional labor to deploy in the environment.
+To acquire high-quality scanning results and avoid motion blur caused by aggressive sensor movement [35], it takes 15 minutes to one hour to finish the sampling of an object, depending on the geometric complexity of the object.
+Generating 3D assets [9, 10, 7, 36] is of great significance because it allows us to learn from limited real-world data with constrained renderings, enabling the creation of more diverse, photorealistic, and scalable 3D content for applications in robotics, computer vision, and beyond.
+The field of 3D reconstruction and generation has seen significant advancements in recent years, driven by the development of neural rendering techniques.
+Neural Radiance Fields (NeRF) [13] pioneered the use of coordinate-based neural networks to represent 3D scenes as continuous volumetric functions, enabling high-quality novel view synthesis.
+However, NeRF’s reliance on dense input views and its computationally intensive training process limit its scalability.
+To address these issues, NeuS [37] introduced a signed distance function (SDF)-based [17, 18] representation, which improves surface reconstruction accuracy by explicitly modeling the geometry of objects.
+This approach achieves better performance in scenarios with sparse input views and complex topologies.
+Building on these advancements, recent works have explored efficient and compact representations for 3D data.
+Among these, the triplane latent representation proposed in [10] is a milestone for 3D generation, where 3D shapes are encoded into three orthogonal feature planes.
+This representation strikes a balance between expressiveness and computational efficiency, enabling faster training and inference while maintaining high fidelity.
+Leveraging the triplane latent representation, researchers have integrated it with generative models to achieve promising results in 3D synthesis.
+GAN-based methods [7, 38] utilize adversarial training to generate realistic 3D shapes by learning the distribution of triplane codes from large-scale datasets.
+On the other hand, DMs [39, 11] have recently gained traction for their ability to generate highly detailed and coherent 3D structures.
+By iteratively denoising triplane codes, diffusion-based approaches achieve superior control over the generation process, enabling applications such as text/image-to-3D.
+Figure 3: Images rendered by our solution without label embedding.
+III Method
+III-A Motivation of Using GANs instead of DMs
 Although diffusion models have recently gained an increasing popularity in 3D object generation, most existing works [20, 11] have only been verified to work under specific category conditions (e.g., chairs [20] or avatars [11]).
@@
 Moreover, given the success of StyleGAN2 [9] in representing NeRFs in latent space, we have chosen latent generation for our pipeline.
-Solution Pipeline.
+III-B Solution Pipeline
 The framework of our solution is depicted in Fig.
@@
 Finally, the attention module produces the 2D synthesis image and a discriminator [10] which only functions during training is used for adversarial training.
-Model Training.
-Since we are simultaneous optimizing NeRFs and SDFs representations, we ensure the generator can initially generate a unit sphere by initializing our model with SDF-pretraining as [18].
+III-C Model Training
+Since we are simultaneously optimizing NeRFs and SDFs representations, we ensure the generator can initially generate a unit sphere by initializing our model with SDF-pretraining as [7].
 Then, the generator is optimized using the Adam optimizer with a learning rate of 0.0025, while the discriminator [10] uses a learning rate of 0.002.
@@
 Additionally, we also use adaptive discriminator augmentation [40] to improve model generalization.
-Overall, the model training objective is formulated as:
+Overall, the model training objective is:
 $\alpha_{0}\mathcal{L}_{path}+\alpha_{1}\mathcal{L}_{e}+\alpha_{2}\mathcal{L}_{v}+\alpha_{3}\mathcal{L}_{sdf},$ (1)
@@
 We set $\alpha_{0}{=}2$, $\alpha_{1}{=}0.05$, $\alpha_{2}{=}0.5$ and $\alpha_{3}{=}0.1$.
-Figure 2: Visualization results of the synthesis images rendered from the NeRFs generated by our solution.
-3 Experiments
-Dataset.
+IV Experiments
+IV-A Dataset
 We train our model from scratch for 3D object generation on the training dataset provided by the ICCV 2023 OmniObject3D Challenge.
 The complete OmniObject3D dataset [1] contains 216 classes from a wide range of daily categories and approximately 6k objects in total where each object has 100 images in different views.
-Efficient Model Training.
+IV-B Efficient Model Training
 Instead of training one model for each class, we train a single model with various types of objects.
-The model takes each image along with its camera pose, focal length and object type.
+The model takes each image along with its camera pose, focal length, and object type.
 Besides, we empirically find that we can effectively train this model with only a few images per object from various classes, so we select the first 8 images of each object for model training.
-This significantly improves the model training efficiency.
-Figure 3: Visualization results of the synthesis images rendered from the NeRFs generated by our solution without additional label embedding.
+This improves the model training efficiency.
+TABLE I: Improvement in terms of FID of our solution after using the additional label embedding.
 Method FID
-$\Delta$= Ours w/ label embedding - -4.8
-Ours w/o label embedding
-Table 1: Improvement in terms of FID of our solution after using the additional label embedding.
-Results.
+NFI (CVPR’23) [7] 62.426
+Ours w/o label embedding 62.335
+Ours w/ label embedding 57.624
+IV-C Results
 Fig.
@@
 3 shows results of our solution without the additional label embedding.
-Comparing Fig.
+By comparing Fig.
 2 with Fig.
-3, we can see that without the additional label embedding, the quality of the synthesis images is significantly worse than the original solution.
-Furthermore, Table.
-1 shows that using the additional label embedding can improve the solution by 4.86 in terms of FID.
-This verifies that the additional label embedding is beneficial for improving model generalization.
-4 Conclusion
-In this technical report, we present a solution for 3D object detection of ICCV 2023 OmniObject3D Challenge.
-The key insight of our solution is to learn effective NeRFs and SDFs Representations with 3D generative adversarial networks for 3D object generation.
-Overall, this solution is one of the final top-3-place solutions in the ICCV 2023 OmniObject3D Challenge.
+3 and the results in Table I, we can see that without the additional label embedding, the quality of the synthesized images is significantly worse than that of the solution with label embedding.
+This verifies that the additional label embedding is beneficial for improving model generalization and performance.
+Furthermore, we compare our method with a state-of-the-art approach proposed in CVPR 2023, nerf-from-image (NFI) [7], as shown in Table I, in terms of FID.
+IV-D Discussion
+Our solution is one of the follow-up works of EG3D [10] that leverages 3D GANs and triplane latent representation.
+On the other hand, the combination of DMs and triplane latent representation [11, 39] has gained significant popularity in recent years.
+In fact, after the announcement of the winners of the ICCV 2023 OmniObject3D Challenge, we found that almost all the competitors used DM-based approaches.
+However, as introduced in Section II and Section III-A, the superiority of DMs lies in their stability during training and control during the generation process (i.e., conditional generation), while the OmniObject3D Challenge is an unconditional generation challenge.
+Thus, it is an interesting topic to explore how well 3D GANs can perform, especially when researchers tend to favor DMs as the preferred solution.
+From this perspective, our work provides some valuable and inspiring results, demonstrating that 3D GANs are still in the game.
+V Conclusion
+In this paper, we present a solution for 3D object generation in the ICCV 2023 OmniObject3D Challenge.
+The key insight of our approach is to learn effective Neural Radiance Fields (NeRFs) and Signed Distance Functions (SDFs) representations using 3D generative adversarial networks (GANs), enabling high-fidelity 3D object synthesis from multiple viewpoints.
+Our experimental results demonstrate that the proposed method effectively generates high-quality images with fine details and accurate geometry.
+We further find that incorporating label embeddings is highly beneficial for improving the model’s learning effectiveness, leading to better generalization and performance.
+Additionally, our approach achieves competitive results compared to a state-of-the-art method proposed at CVPR 2023, further validating its effectiveness.
+Overall, our solution ranks among the top three in the ICCV 2023 OmniObject3D Challenge, highlighting that the 3D GAN-based method is competitive in an era where diffusion models are becoming the dominant approach for 3D generation.